It costs Elmwood, Inc. $73 per unit to manufacture 1,700 units per month of a product that it can sell for $97 each. Alternatively, Elmwood could sell the units at an earlier stage of processing, which would save $34 per unit. Elmwood could sell the simpler product for $64 each. How would selling the simpler product affect Elmwood's profit?

  • Profit would decrease by $56,100.

  • Profit would decrease by $1,700.

  • Profit would increase by $56,100.

  • Profit would increase by $1,700.
Answer #1

Cost price per unit = $73

Selling price per unit = $97

Profit per unit = Selling price per unit-Cost price per unit

= 97-73

= $26

If the product is sold at an earlier stage, cost saving will be $34 per unit.

Hence cost per unit of simpler product = 73-34

= $39

selling price per unit of simpler product = $64

Profit per unit of simpler product = selling price per unit of simpler product-cost per unit of simpler product

= 64-39

= $25

Hence, Profit per unit simpler product is $1 less than if the product is purchased further.

Hence if simpler product is sold, profit would decrease by = 1,700 x 1

= $1,700

Second option is the correct option.

Profit would decrease by $1,700.

Kindly give a positive rating if you are satisfied with this solution and please ask if you have any query.

Thanks